# LLM-SQL-Solver: Can LLMs Determine SQL Equivalence?

### v3

@@ 4.2 Judging Semantic Equivalence
| Judge | Prompt | Noneq | Eq |
| GPT 35 turbo | CounterExample | .8 | .45 |
| Miniature&Mull | 1.0 | .57 | |
| GPT 4 turbo | CounterExample | .6 | .7 |
| Miniature&Mull | 1.0 | .83 |
@@
Recall, two queries can not be semantically equivalent if there exists a counter example (a database instance) such that two queries produce different outputs. We used two different prompting methods Counter Example and Miniature & Mull to compare and contrast the accuracy of LLMs in determining semantic equivalence.
| Judge | Data | Equivalence/Prompt | Easy | Medium | Hard | Extra | Acc |
| - | Spider | - | 248 | 446 | 174 | 166 | - |
| EX | Chat | - | 221 | 331 | 101 | 72 | 70.1% |
| DIN | - | 228 | 382 | 132 | 101 | 81.5% | |
| DAIL | - | 226 | 392 | 132 | 104 | 82.6% | |
| GPT 3.5 turbo | Chat | Semantic/Miniature&Mull | 210 | 293 | 84 | 48 | 61.4% |
| DIN | Semantic/Miniature&Mull | 207 | 307 | 96 | 76 | 66.3% | |
| DAIL | Semantic/Miniature&Mull | 224 | 317 | 85 | 95 | 70.4% | |
[…]
| DIN | Relaxed/Explain&Compare | 223 | 360 | 133 | 114 | 80.3% | |
| DAIL | Relaxed/Explain&Compare | 218 | 371 | 140 | 119 | 82.0% | |
| GPT 4 turbo | Chat | Semantic/ Miniature&Mull | 199 | 244 | 78 | 55 | 55.7% |
| DIN | Semantic/Miniature&Mull | 209 | 308 | 99 | 73 | 66.6% | |
| DAIL | Semantic/Miniature&Mull | 216 | 343 | 114 | 75 | 73.0% | |
[…]
| DIN | Relaxed/Explain&Compare | 209 | 377 | 124 | 106 | 79.0% | |
| DAIL | Relaxed/Explain&Compare | 220 | 382 | 141 | 110 | 82.5% |
As shown in Table 1, we can observe that the Miniature & Mull prompt always provides a higher accuracy than the Counter Example prompt. When using the Counter Example prompt (asking LLMs to determine SQL equivalence and provide a counter example if the SQL queries are not equivalent), GPT 4 turbo failed on a user found bug originally posted on Hackers News lukaseder . This bug contains two SQL queries where one used DISTINCT in selection and the other used UNION ALL to union two sub-queries. Hence, GPT 4 turbo failed to recognize that UNION ALL will not remove duplicate rows (though UNION will). GPT 3.5 turbo failed on the CSE344 exam question which involves different sub-queries. When using Miniature & Mull (asking LLMs to first run a simple example over the given SQL queries and then examining if the example database can be modified to generate a counter example), GPT 3.5 turbo and GPT 4 turbo both successfully identified all the mistakes. Our proposed Miniature & Mull prompt not only increased the accuracy of identifying non-equivalent SQL queries to 100% for both models but also increased the accuracy for identifying equivalent SQL queries by at least 12%. As a result, we believe that letting LLMs execute SQL queries on a simple database instance and then assessing whether a counterexample can be generated by modifying the database is a more effective method to determine SQL semantic equivalence.
4.3 Judging NL2SQL Generation
There are two notions that one can use in judging the equivalence of two SQL queries, i.e., the semantic and relaxed notions. We use the proposed Miniature & Mull and Explain & Compare prompts to evaluate the equivalence of SQL queries following semantic and relaxed notions respectively.
In Table 2, we present the breakdown of the alignment between human preference and the execution match/LLM judgement on the SQL pairs (gold SQL from Spider and generated SQL from Chat, DIN, and DAIL). We observe that the relative ranking among Chat, DIN, and DAIL are preserved for both EX and LLM judgements. Since the generated SQL queries in Chat are from ChatGPT with zero-shot prompting, we expect the SQL queries from DIN and DAIL to achieve better accuracy. Since DAIL applied the self-consistency techniques (resource-intensive Gao et al. (2023)), we expect DAIL to be more likely to produce the best answers. As shown in Table 2, the ranking among Chat, DIN, and DAIL based on execution accuracy, GPT 3.5 turbo judgments, and GPT 4 turbo judgements with varying prompts, all aligned with our expectation.
4.3.1 Agreement between LLMs’ Judgements and Execution Accuracy
For a majority of the questions, LLMs’ judgements and EX agree with each other. In particular, as shown in, Table 3, agreements between LLMs and EX for the semantic equivalence notion have at least 69.6%; similarly, the agreements between LLMs and EX for the relaxed equivalence notion have at least 82.4%. We believe this is an encouraging signal that implies the good quality of the Spider benchmark and LLMs’ judgements. The Spider benchmark contains high quality data points (1000 human hours spent Yu et al. (2018)). Creating these high quality databases, which is necessary for good EX, are expensive and does not scale. Assuming the underlying database is well designed and the EX is highly accurate, then a high agreement between EX and LLMs are obviously desirable.
| Judge | Equiv | Data | Agreement |
| GPT 3.5 turbo | Semantic | Chat | 69.6% |
| DIN | 72.6% | ||
| DAIL | 74.6% | ||
[…]
| DIN | 78.9% | ||
| DAIL | 84.6% | ||
| GPT 3.5 turbo | Relaxed | Chat | 83.8% |
| DIN | 88.4% | ||
| DAIL | 90.0% | ||
[…]
| DIN | 82.4% | ||
| DAIL | 86.1% |
4.3.2 Human Evaluation on Samples
Although, the agreements between execution accuracy and LLM’s judgments are high, there are still lots of questions with disagreements. In this section, we provide a detailed analysis on the sampled questions from the set of questions where EX and LLM’s judgement differs. Since there are hundreds of such questions (different results between EX and LLM) per data source and per GPT models, we sample 10 questions randomly from the SQL query pairs where LLMs and EX disagree.
Equivalence Evaluated Manually To evaluate the equivalence between two SQL queries, we use the human preference approach. Three of the authors manually checked the equivalence of the sampled questions where LLMs and EX disagree. In addition, we have also differentiated the equivalence between the semantic notion and relaxed notion (in essence, the former is objective and the latter is subjective). To ensure the quality of human judgement, the annotations are conducted independently, and we use the majority vote to determine the label for each pair of SQL queries.
| Judge | Equiv | Data | Prefer |
| GPT 3.5 turbo | Semantic | Chat | 2/10 |
| DIN | 4/10 | ||
| DAIL | 5/10 - | ||
[…]
| DIN | 6/10 | ||
| DAIL | 5/10 - | ||
| GPT 3.5 turbo | Relaxed | Chat | 6/10 |
| DIN | 5/10 | ||
| DAIL | 6/10 | ||
[…]
| DIN | 6/10 | ||
| DAIL | 7/10 |
Semantic Equivalence
@@ 4.3.2 Human Evaluation on Samples
When using the semantic equivalence notion, we find GPT 3.5 turbo may not deliver better judgment than EX, but GPT 4 turbo results in better alignments. When EX detects discrepancy in the SQL queries’ outputs, it often indicates these two SQL queries are not semantically equivalent. There are a few rare cases that we observed in which the SQL queries are semantically equivalent but EX determines not equivalent. For example, when the provided database in Spider contains column values that can not be decoded (similar issues are also raised sythello , semantically equivalent queries can be misclassified by EX.
In Table 4, GPT 3.5 turbo aligned with human experts in the range of two to five out of ten. Diving deeper into the questions which GPT 3.5 turbo wrongly classified the SQL queries, we find that there are common reasons for the wrong classifications: Distinct entries vs duplicated entries. GPT 3.5 turbo often failed to consider the difference between selecting unique rows and selecting rows that may be duplicated; Subtle difference in column names. When the gold SQLs and generated SQLs contain only subtle differences in the columns used such as group by id v.s. group by student_id, GPT 3.5 turbo may fail to notice these differences; Hallucination. GPT 3.5 turbo may return wrong results when executing the SQL queries on its own created simple database instance and sometimes when the outputs are indeed identical, it may still state that the outputs are different. This hallucination snowballing effect Zhang et al. (2023a) is one of the top reason why GPT 3.5 turbo performs worse compared to GPT 4 turbo.
When GPT 4 turbo is used as the judge, it constantly produced more accurate or comparable results to EX, with at least 5 out of 10 questions for DAIL and at most 10 out of 10 questions for Chat. One interesting observation is that GPT 4 turbo correctly identified the difference between order by asc/desc limit 1 v.s. where col = select min/max(col). The former will always return one row, as limit 1 is used. The latter, however, may return many rows when these rows share the same min or max value. EX often failed to capture this difference. Although GPT 4 turbo’s judgment is much better, GPT 4 turbo may still miss some minor differences such as similar but different column names (such as song_name v.s. name), overlook the fact that extra columns are selected, and also sometimes hallucinates. As shown in Figure 2, we explored how well do LLMs execute and reason about the SQL queries on their own generated simple database. GPT 4 turbo has much higher correct operation rates which leads to much better responses compared with the weaker model. One common phenomena that lead to false answers in EX is due to the return of the empty tuple. When the underlying database does not contain tuples that match the predicate, gold SQLs and generated SQLs will always match, irrespective of how different they are. Another interesting observation raised from a few comparisons is about the case sensitivity. OpenAI GPT models consider databases are by default case insensitive; Spider benchmark, on the other hand, uses the SQLite databases which are case sensitive by default. However, many popular database are by default case insensitive such as SQL Server Mistry and Misner (2014) and MySQL DuBois (2013). Therefore, we operate under the assumption that the underlying database instance is case-insensitive when evaluating semantic equivalence.
@@
Relaxed Equivalence. When using the Explain & Compare prompt for determining the relaxed equivalence, LLMs are consistently better aligned with human judgments than EX, as shown in Table 4. The relaxed notion aims to illustrate the degree to which humans perceive two queries as being similar in their meaning, and mimics the opinions of users. The positive results on the alignment between LLMs’ judgments with human preferences indicate optimizing accuracy for LLMs’ judgments is a better optimization target than EX.
The main reasons for why LLMs’ judgements are better aligned with human preference than EX is that LLMs can zoom into the query logic and determine whether the two SQL queries share the same logic and contain the same truth. Consider the following example, order by col asc/desc limit 1 v.s. where col = select min/max(col). LLMs can correctly identify that these two queries are selecting the rows with minimal or maximal value. Since EX strictly checks the outputs, it may sometimes return equivalent and other times return not equivalent. One interesting observation we found is about null values. GPT often consider COUNT(*), COUNT(primary key), and COUNT(DISTINCT primary key) to be different. The reason is that GPT models consider that null values can be inserted into a column with primary key constraint. In fact, the SQLite databases, by default, allow null values to be inserted to primary key column. However, in many database primary key constraints explicitly states that null values are not allowed Oracle . Based on the relaxed equivalence definition, we consider records in the primary columns to be not null. Although LLMs correctly consider that having null values in the primary key column may lead to different results (note that COUNT(*) considers null values as unique entries, where as COUNT(primary key) ignores rows with null values), we consider these not as a significant issue.
5 Discussion and Limitation
In this discussion section, we would like to highlight the broader implications and potential integration of our proposed LLM-SQL-Solver. During the SQL rewrite phase of the query optimization engine, the input SQL query is transformed into a new SQL query that can be optimized more easily. The key challenge in query rewriting is to determine if the input query and the rewritten query are indeed semantic equivalent. The LLM-SQL-Solver can be directly used as a tool, during the rewriting phase, to check semantic equivalence and ensure bugs (e.g., COUNT bug Ganski and Wong (1987)) do not occur in the data system. Furthermore, the integration of the LLM-SQL-Solver as a text-to-SQL evaluation metrics using the relaxed equivalence notion could improve the overall quality of user experience and enable more human-like capabilities.

### v4

@@ 4.2 Judging Semantic Equivalence
| Judge | Prompt | Semantic Eq. | Semantic Ineq. | ||
| 0-Shot | 5-Shot | 0-Shot | 5-Shot | ||
| ChatGPT | CoT | 53% | 52% | 63% | 46% |
| M&M | 85% (+32%) | 87% (+35%) | 67% (+4%) | 47% (+1%) | |
| GPT-4o | CoT | 81% | 83% | 85% | 83% |
| M&M | 84% (+3%) | 89% (+6%) | 86% (+1%) | 85% (+2%) | |
@@
Recall that two queries cannot be semantically equivalent if there exists a counter example (a database instance) such that two queries produce different outputs. We used two different prompting methods CoT and Miniature & Mull to compare and contrast the accuracy of LLMs in determining semantic equivalence.
As shown in Table 1, we can observe that the Miniature & Mull prompt always provides a higher accuracy than the CoT prompt. For determining the semantic equivalent queries (Calcite test suite), Miniature & Mull prompt using GPT-4o achieved around 90% accuracy. For determining the semantic inequivalent queries (DAIL-SQL results over Spider Benchmark), Miniature & Mull prompt using GPT-4o achieved around 86% accuracy. We can also observe that when using the weaker GPT-3.5-Turbo in judging semantically equivalent queries, Miniature & Mull outperform CoT prompt (asking LLMs to determine SQL equivalence and think step by step) by more than 30%. However, when judging the inequivalent SQL queries, we find that the GPT-3.5-Turbo significantly under-perform than the GPT-4o Turbo model.
Analyzing responses for identifying SQL-inequivalent queries, we observed a few reasons for GPT-3.5-Turbo’s low accuracy:
1.
Distinct entries vs duplicated entries. GPT-3.5-Turbo often failed to consider the difference between selecting unique rows and selecting rows that may be duplicated.
2.
Case-Sensitivity Ignored: GPT-3.5-Turbo often disregards the fact that the database is case-sensitive. Althought we have explicitly noted the database is string case-sensitive in the prompt, GPT-3.5-Turbo often fails to strictly adhere to the case-sensitivity requirement, resulting in inaccuracies when comparing queries. This observation aligns with findings in prior work Wen et al. (2024), which highlight GPT-3.5-Turbo’s challenges in reliably following instructions compared to stronger models such as GPT-4o.
3.
Overlooked Subtle Differences: GPT-3.5-Turbo may also fail to detect critical differences in column and table names. For instance, it might incorrectly treat distinct column names as equivalent or overlook the inclusion of additional columns or elements in a query. A specific example includes comparing AVG(capacity) with avg_capacity, where the former performs an average operation on the capacity column, while the latter refers to an entirely different column. Similarly, in cases like GROUP BY id versus GROUP BY student_id, GPT-3.5-Turbo may fail to recognize these distinctions, potentially leading to incorrect judgments.
4.
Hallucination. In the Miniature & Mull prompt, LLMs are instructed to execute both SQL queries on a simple database they create and then repeat the process on a modified version of the database to compare outputs. GPT-3.5-Turbo often returns incorrect results when executing SQL queries on its self-generated database instances. Moreover, even when the outputs are identical, it may mistakenly assert that the results differ. This hallucination and the resulting snowballing effect Zhang et al. (2023a) significantly contribute to GPT-3.5-Turbo’s poorer performance compared to GPT-4o.
When GPT-4o is used to determine SQL semantic equivalence, it consistently produces more accurate and reliable decisions for both equivalent and inequivalent queries. A notable observation is that GPT-4o reliably identifies the distinction between ORDER BY ASC/DESC LIMIT 1 and WHERE col = (SELECT MIN/MAX(col)). The former always returns a single row due to the LIMIT 1 clause, whereas the latter may return multiple rows if several share the same minimum or maximum value. This critical difference is often overlooked by GPT-3.5-Turbo. When using the CoT (Chain-of-Thought) prompt, GPT-4o achieves an accuracy of 82% – 85%, and with the Miniature & Mull prompt, it improves further to 85% – 90%.
Furthermore, we observe that, in general, adding few-shot demonstrations improves accuracy. However, for semantically inequivalent queries, incorporating the five static examples results in a significant drop in accuracy for both the CoT and Miniature & Mull prompts when using GPT-3.5-Turbo. The same few shot approach leads to only a slight accuracy decline when using GPT-4o. This discrepancy underscores the superior in-context learning capabilities of GPT-4o compared to GPT-3.5-Turbo. GPT-4o demonstrates a greater ability to interpret and generalize from the provided examples, enabling it to more effectively account for nuanced differences in semantic meaning. Conversely, GPT-3.5-Turbo struggles with leveraging the few-shot examples for semantically inequivalent queries, likely due to limitations in its contextual reasoning and pattern recognition abilities. This results in a diminished capacity to differentiate between queries with subtle semantic differences. Moreover, we believe that determining SQL semantic equivalence is an inherently challenging task for LLMs. It requires an understanding of both the syntax and the underlying logic of the queries. While few-shot learning provides valuable context, it may not always suffice to overcome the complexities involved, particularly for cases where the differences between queries are subtle or deeply tied to specific database schema (e.g., Primary Key and Foreign Key relationships).
Despite the aforementioned challenges in determining SQL semantic equivalence, our proposed Miniature & Mull prompt always achieve higher accuracy than using Cot prompt. This demonstrates the effectiveness of our approach in addressing some of the inherent difficulties of the task. As a result, we believe that letting LLMs execute SQL queries on a simple database instance and then assessing whether a counterexample can be generated by modifying the database is a more effective method to determine SQL semantic equivalence.
@@
To conclude this subsection, we also note that there are a few very rare cases that we observed in which the SQL queries are semantically equivalent but execution match determines not equivalent. For example, when the provided database in Spider contains column values that can not be decoded (similar issues are also raised sythello , semantically equivalent queries can be misclassified by execution match but very rare.
| Judge | Data | Easy | Medium | Hard | Extra | Overall Acc | Rank |
| - | Spider Benchmark | 248 | 446 | 174 | 166 | - | - |
| EX | Chat-SQL | 221/248 | 331/446 | 101/174 | 72/166 | 70.1% | 3 |
| DIN-SQL | 228/248 | 382/446 | 132/174 | 101/166 | 81.5% | 2 | |
| DAIL-SQL | 226/248 | 392/446 | 132/174 | 104/166 | 82.6% | 1 | |
| ChatGPT | Chat-SQL | 223/248 | 397/446 | 149/174 | 128/166 | 86.8% | 3 |
| DIN-SQL | 228/248 | 391/446 | 153/174 | 133/166 | 87.5% | 2 | |
| DAIL-SQL | 235/248 | 407/446 | 158/174 | 142/166 | 91.1% | 1 | |
| GPT-4o | Chat-SQL | 226/248 | 328/446 | 129/174 | 104/166 | 76.1% | 3 |
| DIN-SQL | 231/248 | 376/446 | 150/174 | 126/166 | 85.4% | 2 | |
| DAIL-SQL | 228/248 | 397/446 | 153/174 | 129/166 | 87.7% | 1 |
4.3 Judging Text-to-SQL Generation
In this subsection, we investigate whether large language models (LLMs) can effectively assess the quality of text-to-SQL generations. Specifically, we compare our proposed Explain & Compare method with the widely used execution match approach. Additionally, we examine which evaluation metric (i.e., LLM judgments or execution match) aligns more closely with user opinions, providing deeper insights into the effectiveness of these approaches.
We first introduce three state-of-the art text-to-SQL algorithms for the Spider benchmark.
DIN-SQL Pourreza and Rafiei (2023) ranks the fourth in the leader board and it decompose the text-to-SQL translation problem into sub tasks such as schema linking and classifying the level of difficulties. DIN-SQL uses the GPT-4 model with few-shot prompting.
Chat-SQL Huang et al. (2023) directly generate queries using GPT-3.5-Turbo with the zero-shot basic prompt.
In Table 2, we present the breakdown of the alignment between human preference and the execution match and the LLM judgment using Explain & Compare prompt on the SQL pairs (gold SQL from Spider and generated SQL from Chat, DIN, and DAIL).
First, we observe that the relative ranking among Chat-SQL, DIN-SQL, and DAIL-SQL remains consistent when transitioning from execution match accuracy to LLM-based evaluations. Given that Chat-SQL relies on SQL queries generated by the weaker model, GPT-3.5-Turbo, using zero-shot prompting, it is expected that the SQL queries produced by DIN-SQL and DAIL-SQL would demonstrate higher accuracy. Furthermore, DAIL-SQL stands out by extensively exploring various prompting techniques for text-to-SQL translation, employing the most effective combination of strategies, and incorporating resource-intensive self-consistency methods. As a result, DAIL-SQL is more likely to generate the most accurate responses. As shown in Table 2, the ranking among Chat-SQL, DIN-SQL, and DAIL-SQL based on execution accuracy, GPT-3.5-Turbo judgments, and GPT-4o judgments with our proposed Explain & Compare, all aligned with our expectation.
Additionally, when we examine the accuracy across different difficulty levels (i.e., easy, medium, hard, and extra) in the Spider benchmark, it becomes evident that the primary discrepancies between the execution metric and the LLMs’ judgment metric arise from questions classified as hard and extra difficulty. LLMs’ judgments, based on the pragmatic equivalence, are less strict than the execution match metric. Consider the following example, order by col asc/desc limit 1 v.s. where col = select min/max(col). LLMs using Explain & Compare prompt are able to identify the pragmatic equivalence between these two queries (i.e., identifying the top/bottom results based on col) despite their syntactic differences. However, since execution match strictly checks the outputs, it may sometimes return equivalent and other times return not equivalent depending on the underlying database (i.e., whether there are duplicate records sharing the same maximum or minimal value).
| Judge | Data | Agreement |
| GPT-3.5-Turbo | Chat-SQL | 75.8% |
| DIN-SQL | 83.4% | |
| DAIL-SQL | 85.5% | |
| GPT-4o | Chat-SQL | 82.6% |
| DIN-SQL | 87.2% | |
| DAIL-SQL | 89.5% |
4.3.1 Agreement between LLMs’ Judgments and Execution Accuracy
For the majority of questions, the judgments made by large language models (LLMs) align well with execution-based (EX) accuracy. Specifically, as shown in Table 3, the agreement rates between LLMs’ judgments (based on pragmatic equivalence) and execution accuracy range from 76% to 90%. This strong alignment is an encouraging signal, reflecting both the quality of the Spider benchmark and the reliability of LLMs’ judgments. The Spider benchmark is renowned for its high-quality databases, developed with significant human effort—approximately 1,000 hours invested in their creation Yu et al. (2018). This meticulous design ensures the databases are robust, enabling execution-based metrics to produce reliable results. The benchmark’s reputation has been widely celebrated within the community for setting a standard in text-to-SQL evaluation.
However, creating such high-quality databases is resource-intensive and does not easily scale, limiting the broader applicability of execution-based metrics. In this context, the high agreement between execution accuracy and LLM-based judgments is particularly desirable. It suggests that LLMs, when guided by pragmatic equivalence, can provide a scalable and reliable alternative for evaluating query accuracy, reducing the dependence on expensive database creation while maintaining strong evaluation standards.
4.3.2 Alignment with Human Preference
While LLM judgments and execution-based metrics agree on the majority of questions, there remain many instances where their evaluations diverge. In this subsection, we analyze which metric, LLM judgments or execution accuracy, better aligns with human preferences. By examining these discrepancies, we aim to uncover the strengths and weaknesses of each evaluation approach in accurately assessing the quality of text-to-SQL systems.
| Judge | Few Shot | Human Prefer |
| Execution Match | - | 40.0% |
| GPT-3.5-Turbo | 0 Shot | 72.9% |
| 5 Shot | 74.3% | |
| GPT-4o | 0 Shot | 68.6% |
| 5 Shot | 78.6% |
Human Preferences To evaluate the pragmatic equivalence between two SQL queries in text-to-SQL service, we use the human preference approach. Higher alignment to human preference indicates better judgment. At least three experts in SQL have manually checked the equivalence of the 70 selected queries from the state-of-the-art text-to-SQL algorithms Pourreza and Rafiei (2023); Gao et al. (2023); Huang et al. (2023). These 70 questions include 9 easy questions, 27 medium questions, 14 hard questions, and 20 extra-hard questions. They were randomly selected from cases where an LLM’s judgment disagreed with the execution match metric. To ensure the quality of human judgment, the annotations are conducted independently, and we use the majority vote to determine the final label for each pair of SQL queries.
@@ 4.3.2 Alignment with Human Preference
In Table 4, we find that execution match only aligned 40% of the time with human preferences. When using the Explain & Compare prompt for determining the pragmatic equivalence, LLMs are consistently better aligned with human preference than execution match, as shown in Table 4. The pragmatic notion aims to illustrate the degree to which humans perceive two queries as being similar in their meaning and intent, mimicking the opinions of text-to-SQL users. By reasoning into the query logic and determine whether the two SQL queries share the same logic and have the same intent, LLMs’ judgments become better aligned with human preference than execution accuracy. Moreover, when five static examples are included in the prompt, GPT-3.5-Turbo aligned with human preference 74.3% and GPT-4o aligned with human preference 78.6%. The positive results on the alignment between LLMs’ judgments with human preferences indicate optimizing accuracy for LLMs’ judgments is a better optimization target than EX.
Diving deeper into the questions where LLMs failed to align with human preferences, we find that these questions often lack unanimous agreement among experts, as the definition of pragmatic equivalence can be highly subjective. For instance, consider the following two SQL queries:
The queries differ in their filtering criteria: the first uses the airport name to filter for ’Alton’, while the second uses the airport code to filter for ’Alton’. Among experts, two considered these queries pragmatically inequivalent, as the filtering criteria are based on different attributes, which could yield totally different output. However, one expert judged them to be pragmatically equivalent, as they have the same intent of finding location based on predicate ’Alton’. LLMs, failing to converge to the majority opinion, interpreted these queries as having the same meaning, possibly oversimplifying the nuanced distinctions in their semantic intent. An exciting avenue for future research is to develop customizable text-to-SQL evaluation metrics tailored to individual user preferences, leveraging in-context learning to adapt the system’s understanding and judgments based on specific user requirements or domain contexts.
5 Discussion and Future Works